40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 14 23:14:24 2018, aaron, Update, OMC, OMC scanning/aligning script 
    Reply  Tue Nov 20 20:33:11 2018, aaron, Update, OMC, OMC scanning/aligning script pathwaysketch.pdfIMG_0094.JPGIMG_0097.JPG
       Reply  Mon Nov 26 15:43:16 2018, aaron, Update, OMC, OMC scanning/aligning script c1omc.png
          Reply  Thu Dec 6 11:16:28 2018, aaron, Update, OMC, OMC channels 
             Reply  Mon Dec 10 12:29:05 2018, aaron, Update, OMC, OMC channels 52DE723A-02A4-4C62-879B-7B0070AE8A00.jpeg545E5512-D003-408B-9F00-55F985966A16.jpegDFF34976-CC49-4E4F-BFD1-A197E2072A32.jpeg
                Reply  Mon Dec 10 19:47:06 2018, aaron, Update, OMC, OMC channels 
                   Reply  Tue Dec 11 13:42:44 2018, Koji, Update, OMC, OMC channels 
                   Reply  Tue Dec 11 13:48:04 2018, aaron, Update, OMC, OMC channels 
                      Reply  Thu Dec 13 22:24:21 2018, aaron, Update, OMC, OMC channels 
          Reply  Mon Dec 10 12:11:28 2018, aaron, Update, OMC, Aligning the OMC B9DCF55F-1355-410C-8A29-EE45D43A56A4.jpeg
             Reply  Tue Dec 11 14:24:18 2018, aaron, Update, OMC, Aligning the OMC 
                Reply  Tue Dec 11 22:50:07 2018, aaron, Update, OMC, Aligning the OMC 6x
                   Reply  Thu Dec 13 22:36:42 2018, aaron, Update, OMC, Aligning the OMC 
                      Reply  Fri Dec 14 13:02:24 2018, aaron, Update, OMC, Aligning the OMC 
                      Reply  Fri Dec 14 13:06:12 2018, aaron, Update, OMC, Aligning the OMC 
                         Reply  Mon Dec 17 20:45:40 2018, aaron, Update, OMC, Aligning the OMC IMG_5922.JPGIMG_5921.JPG
                            Reply  Tue Dec 18 18:13:32 2018, aaron, Update, OMC, OMC L HV Piezo driver tests (again) IMG_4576_copy.jpg
                               Reply  Wed Dec 19 00:12:46 2018, gautam, Update, OMC, 40m OMC DCC node 
Message ID: 14363     Entry time: Mon Dec 17 20:45:40 2018     In reply to: 14358     Reply to this: 14365
Author: aaron 
Type: Update 
Category: OMC 
Subject: Aligning the OMC 

[gautam, aaron]

We did work in the OMC chamber today to get the OMC aligned. Aaron was in the clean suit while Gautam steered in-air optics. We modified the aux input steering optics and the final two OMC steering optics (between OMMT and OMC), but did not modify any of the AS path optics.

I had already aligned AUX approximately into the AS port from the AP table. With the OMC N door open, we aligned the aux beam first to OM6, then to OMPO, then OM5. OM5 was the last optic in the OMC chamber that we could align to.

From there, Gautam found the aux beam clipping on a few optics on its way to SR4 using the IR viewer. Once we were approximately hitting SR4, we got a return beam in the OMC chamber, which we were able to coalign with the input aux beam.

We had already done the alignment of SR5 into the OMC during the last vent, so we immediately had a refl off of the OMC, which we aligned onto a PD520 from the PSL table (larger aperture than the previous PD, which anyway needed a macroscopic adjustment to catch the refl beam).

Next, we removed the OMC cover, wrapped it in foil, and placed it in the makeshift clean room near the Y end. The screws remain in a foil bucket in the OMC chamber. With the cover off, Aaron moved the OMC input steering mirrors to align the beam in the OMC. We measured ~2.4mW in the OMC refl beam, which means about 240uW is transmitted into the OMC. Aaron thinks the beam overlaps itself after one round trip in the cavity, but that the entire plane may be too low in pitch, so more alignment may be needed here.

With the beam approximately aligned into the OMC, we energized the OMC-L piezo driver with 200V, and applied a ~0.03Hz triangle wave on the OMC diff input (pins 2-7). We monitor the REFL PD, piezo mon, function generator signal, and one of the trans PDs. We noticed that the PZT mon shows the driver saturating before the function generator reaches its full +-10V, which is something to investigate.

We saw what could have been regular dips in the REFL PD signal, but realized that with an unkown level of mode matching, it will be hard to tell whether the light becomes resonant with the DC signal. Gautam has suggested coaligning the aux and PSL beams, then observing the PDH signal from the PSL beam as the OMC sweeps through resonance, while turning aux back on anytime we try to make adjustments to the alignment of the OMC (so I can see the beam in the cavity).

I'll think through the plan in some more detail and we will try to have the OMC locked tomorrow.


gautam:

  1. All references to SR4 etc actually refer to OM4 etc.
  2. For this experiment, we are using the prompt reflection of the AUX beam from the HR surface of the SRM, so as to get maximum light into the cavity.
  3. For 2.4 mW incident on OMC1 (actually we measured the REFL beam on the AS table), we expect ~24 uW inside the cavity, which isn't a lot but still was visible on the card.
  4. After this work, I checked the IMC alignment - it was still easily able to lock to a TEM00 mode, but the transmission was ~half (i.e. ~600cts) of what I am used to it being in low power mode (~1300 cts). I didn't align the cavity to the input beam, as I think in this case, the right thing to do is to align the input beam to the IMC cavity axis.
Attachment 1: IMG_5922.JPG  4.472 MB  | Hide | Hide all
IMG_5922.JPG
Attachment 2: IMG_5921.JPG  3.956 MB  | Hide | Hide all
IMG_5921.JPG
ELOG V3.1.3-